OpenStudy (anonymous):

-6|5-2y|=-9 I don't just want the answer I want the steps so I can fully understand it.

OpenStudy (lynfran):

-6|5-2y|=-9 1st step divide both sides by -6....wat do we get ?

OpenStudy (anonymous):

Do you have the answer now?

OpenStudy (anonymous):

|5-2y| =1\frac{ 1 }{ 2 } after dividing

OpenStudy (lynfran):

|5-2y|=3/2 next step.. we split the absolute value into 2...having 1 positive and 1 negative ..like this (5-2y)=3/2 and (5-2y)=-3/2 we now can solve the 2 equation for y..